A 30.0g sample of krypton has a temperature of 20Degree celcius at 555Hg what is the volume in mL of krypton gas
1
Expert's answer
2016-10-29T10:46:07-0400
n (Kr) = m/M = 30 / 83.8 = 0.358 (mol)
V (Kr)stp = n*Vm = 0.358 *22.4 = 8 (l)
pV/T = p1V1/T1
V1 = pVT1/p1T = 760 * 8 * 393 / 555 * 373 = 11.54 (l)

Answer
V (Kr) = 11540 mL